OpenStudy (ladiesman217):

MEDAL AND FANS AWARDED help please i dont understand this? The equations only has 2 variables and i thought a quadratic equation requires 3 variables? Use the quadratic formula to solve -6x^2 + 6x = 0

OpenStudy (cwrw238):

terms you mean not variables. there is only one variable in your equation thats x

OpenStudy (cwrw238):

a quadratic equation can have only 2 terms -6x^2 + 6x = 0 here a = -6, b = 6 and c = 0 so plug these values in to the formula and work out the roots

OpenStudy (ladiesman217):

how do you do that though?

OpenStudy (cwrw238):

x = [-b +/- sqert(b^2 -4ac) ] / 2a =[ -6 - sqrt(36 - 4*0*-6) [/ -12 =[ -6 - sqrt36] / 2 = -12 / 2 = -6 and the other value of x is [- 6 + 6] / 2 = 0